OpenStudy (anonymous):

Use the substitution method to solve the following system of equations. 4x – y = 3 7x – 9y = –2

OpenStudy (anonymous):

(1, –3) (6, –3) (1, 1) (6, 1)

OpenStudy (paxpolaris):

from eq 1 you get :y=4x-3

OpenStudy (paxpolaris):

substitute that for y in eq.2 and solve

OpenStudy (paxpolaris):

\[7x-9\left( 4x-3 \right)=-2\]

OpenStudy (anonymous):

i came outwith a solutuion of (1, 1)????

OpenStudy (paxpolaris):

yep

OpenStudy (anonymous):

thank you :)

OpenStudy (anonymous):

first thing you want to do is solve for either x or y. In the first equation, we can solve for y by subtracting 4x by both sides. -y = 3 - 4x remember our variable can't be negative, so multiply everything by -1 y = -3 + 4x we have now solved for y. Now, we will plug in what is equal to y in the second equation where you see a y. in 7x - 9y = -2 you see a y right by 9, so let's plug in -3+4x into it, because that is what y equals 7x -9(-3+4x) = -2 use the distributive property 7x + 27 -36x = -2 combine like terms -31x + 27 = -2 subtract 27 from both sides -29x = -29 x = 1 Now that we know what x equals, let's plug it back into the equation y = - 3 + 4 (1) y = 3 + 4 y = 1

OpenStudy (anonymous):

yeah im late, no i dont care